Mr. Smith is building a large, circular driveway for his plantation home which has an outside diameter of 220 ft, and its soon-to-be concrete surface will be 10 feet wide (inside diameter is 200 ft), with concrete thickness being 6 inches. What is it's surface area in square feet? Let pi=3.14.
Answer by mananth(16946)

Area of outer circle - area of inner circle = area of driveway
outer radius =220/2 = 110
inner radius = 200/2= 100
pi*r1^2-pi*r2^2
pi*(r1^2-r2^2)
pi*(110^2-100^2)
pi*(12100-10000)
pi*2100
=6594 sq.ft